OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (8, 12) and has a slope of −2?

OpenStudy (anonymous):

@perl @iGreen @abswag13 @logancoop @lsherron @tennisdude210

OpenStudy (rainbowcupcake7997):

y=mx+b You have y and x and b is the slope. so first fill that in..

OpenStudy (anonymous):

y = -2x + b and thats how far i have got

OpenStudy (anonymous):

plz help asap i need it

OpenStudy (rainbowcupcake7997):

y=12 x=8 12=-2(8)+b

OpenStudy (rainbowcupcake7997):

So now to find b which is y-intercept. Just turn the problem around and have b alone on one side and then solve.

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

i have calculator with me :P

OpenStudy (rainbowcupcake7997):

then just right it in slop-intercept form. y=mx+b

OpenStudy (rainbowcupcake7997):

wait how'd you get -4?

OpenStudy (anonymous):

brb

OpenStudy (rainbowcupcake7997):

12=-2(8)+b add 2 to both sides 14=(8)+b subtract 8 from both sides 6=b

OpenStudy (anonymous):

hi @RainbowBubbles01